# ActionReasoningBench: Reasoning about Actions with and without Ramification Constraints

### v2

@@ 5.1 Ramifications Results
Ac Seq. Ques Cat GPT-4o o1-preview 1 Fluent Trk ${00.00}_{00.00}$ ${00.00}_{00.00}$ State Trk ${00.00}_{00.00}$ ${20.00}_{17.89}$ Effects of Ac ${00.00}_{00.00}$ ${40.00}_{21.91}$ Average ${00.00}_{00.00}$ ${25.00}_{12.49}$ 10 Fluent Trk ${00.00}_{00.00}$ ${00.00}_{00.00}$ State Trk ${00.00}_{00.00}$ ${33.33}_{19.24}$ Effects of Ac ${00.00}_{00.00}$ ${14.28}_{13.22}$ Average ${00.00}_{00.00}$ ${23.07}_{11.68}$ 19 Fluent Trk ${00.00}_{00.00}$ ${33.33}_{27.21}$ State Trk ${00.00}_{00.00}$ ${00.00}_{00.00}$ Effects of Ac ${00.00}_{00.00}$ ${00.00}_{00.00}$ Average ${00.00}_{0.00}$ ${07.69}_{07.38}$
Table 5: Performance comparison of GPT-4o and o1-preview on the free-answer subset of the benchmark, evaluated with the ramifications constraints using the zero-shot-CoT. The results are categorized up by the action-sequence lengths and question categories.
As discussed in Section 4, the performance on ramification fluents is evaluated for two LLMs: GPT-4o, the highest-performing LLM, and o1-preview, the most recent state-of-the-art LLM. Table 5 presents the performance of both models when ramification constraints are introduced. Further examples of the model responses to ramification-related questions can be found in the Appendix I
@@ 6 Conclusion
Additionally, we explored the indirect effects of actions, known as ramifications, where even state-of-the-art models show considerable limitations. For example, GPT-4o could not solve any questions involving ramifications, and the o1-preview model achieved a low score of 18.4%. These results underscore the pressing need for further research and advancements in RAC reasoning, especially in addressing indirect effects and more advanced reasoning tasks. Improving LLMs’ capabilities in these areas is crucial for enhancing their performance in dynamic and interactive environments. Our findings reinforce the challenges LLMs face in RAC tasks, which, despite their polynomial complexity in cases without ramifications, serve as prerequisites for more advanced planning problems that are often NP-complete or beyond. Improving LLM performance in RAC is a pivotal step toward overcoming broader planning challenges.
7 Limitations & Future Work
@@
    Although there are more complex types of RAC, exploring those remains beyond the scope of this work and is left as a direction for future research.
@@
Appendix A Describing an Instance from ActionReasoningBench
@@
Appendix B Few Shot 3 Results
@@
Appendix C Data Verification
@@
Appendix D Fine-tuning Details
In this section, we describe the fine-tuning performed on the training split of ActionReasoningBench described in section 3.5. We fine-tuned Llama-3.1-8B separately for binary (true/false) and free answer questions, using 6 epochs for the former and 18 epochs for the latter. The AdamW optimizer was used, with a batch size of 4 and gradient accumulation steps set to 8 for both of the training setups. Due to the available compute resources, we were limited to a maximum context length of 4096 tokens. This leaves us roughly with 27k samples for binary answers and 14.4k samples for free answer. Tables 9, 10 and 11 show the statistics of the training set that we used to train the models.
@@ Appendix D Fine-tuning Details
Appendix E Free Answers Evaluation Details
@@
Appendix F Domains in ActionReasoningBench
@@ F.8 VisitAll
Appendix G Planning Description and Tools
@@
G.1 Planning Domain Definition Language (PDDL)

### v3

@@ 5.1 Ramifications Results
As discussed in Section 4, the performance on ramification fluents is evaluated for two LLMs: GPT-4o, the highest-performing LLM, and o1-preview, the most recent state-of-the-art LLM. Table 4 presents the performance of both models when ramification constraints are introduced. Further examples of the model responses to ramification-related questions can be found in the Appendix J.
Action Sequence Question Categories Free Answer Binary Questions GPT-4o o1-preview GPT-4o o1-preview 1 Fluent Tracking ${00.00}_{00.00}$ ${00.00}_{00.00}$ ${71.43}_{17.10}$ ${100.00}_{00.00}$ State Tracking ${00.00}_{00.00}$ ${20.00}_{17.89}$ ${100.00}_{00.00}$ ${100.00}_{00.00}$ Effects of Actions ${00.00}_{00.00}$ ${40.00}_{21.91}$ ${71.43}_{17.10}$ ${57.14}_{18.70}$ Average ${00.00}_{00.00}$ ${25.00}_{12.49}$ ${80.95}_{08.57}$ ${85.71}_{07.60}$ 10 Fluent Tracking ${00.00}_{00.00}$ ${00.00}_{00.00}$ ${57.14}_{18.70}$ ${57.14}_{18.70}$ State Tracking ${00.00}_{00.00}$ ${33.33}_{19.24}$ ${42.86}_{18.70}$ Effects of Actions ${00.00}_{00.00}$ ${14.28}_{13.22}$ ${71.43}_{17.10}$ ${100.00}_{00.00}$ Average ${00.00}_{00.00}$ ${23.07}_{11.68}$ ${57.14}_{10.80}$ ${78.57}_{10.90}$ 19 Fluent Tracking ${00.00}_{00.00}$ ${33.33}_{27.21}$ ${42.86}_{18.70}$ ${57.14}_{18.70}$ State Tracking ${00.00}_{00.00}$ ${00.00}_{00.00}$ ${57.14}_{18.70}$ Effects of Actions ${00.00}_{00.00}$ ${00.00}_{00.00}$ ${71.43}_{17.10}$ ${85.71}_{13.20}$ Average ${00.00}_{00.00}$ ${07.69}_{07.38}$ ${57.14}_{10.80}$ ${71.43}_{12.10}$
Table 4: Performance comparison of GPT-4o and o1-preview on both the free-answer and binary question subset of the benchmark, evaluated with the ramifications constraints using the zero-shot-CoT. The results are categorized up by the action-sequence lengths and question categories with a “-” indicating no response (due to longer prompts). We have provided some responses in App. J.
@@ 6 Conclusion
Additionally, we explored the indirect effects of actions, known as ramifications, where even state-of-the-art models show considerable limitations. For example, GPT-4o could not solve any questions involving ramifications, and the o1-preview model achieved a low score of 18.4%. These results underscore the pressing need for further research and advancements in RAC reasoning, especially in addressing indirect effects and more advanced reasoning tasks.
Acknowledgement
We thank the anonymous reviewers for their constructive suggestions.
We extend our gratitude to the Research Computing (RC), and Enterprise Technology at ASU for providing computing resources, and access to the ChatGPT enterprise version for experiments.
We acknowledge support by a 2023 Spring Amazon Research Award (ARA). This material is also based upon work supported by the Engineering Research and Development Center - Information Technology Laboratory (ERDC-ITL) under Contract No. W912HZ24C0022.
Son Tran acknowledges the partial support from NSF grant #1914635 during the time working on this paper.
Appendix A Limitations & Future Work
@@
    Although there are more complex types of RAC, including incorporating more reasoning types, exploring those remains beyond the scope of this work and is left as a direction for future research.
@@
Appendix B Describing an Instance from ActionReasoningBench
@@
Appendix C Additional Results
C.1 Zero-shot-CoT Binary Result
Table 5 shows the results on GPT-4o, Llama-3.1-8B-Instruct, and Llama-3.1-70B-Instruct on zero-shot-CoT prompting.
C.2 Few-shot-3 Results
@@
C.3 Results by Domains
Act. Seq. Ques Categories Blocksworld Depots Driverlog Grippers Mystery Satellite Spanner Visitall 1 Fluent Tracking ${100.0}_{0.0}$ ${92.31}_{7.69}$ ${69.23}_{13.32}$ ${95.24}_{4.76}$ ${100.0}_{0.0}$ ${88.24}_{8.05}$ ${88.89}_{7.62}$ ${100.0}_{0.0}$ State Tracking ${83.33}_{9.04}$ ${93.33}_{6.67}$ ${66.67}_{16.67}$ ${93.75}_{6.25}$ ${85.71}_{9.71}$ ${61.54}_{14.04}$ ${62.5}_{12.5}$ ${100.0}_{0.0}$ Action Executability ${71.43}_{10.1}$ ${76.47}_{10.6}$ ${76.47}_{10.6}$ ${78.95}_{9.61}$ ${66.67}_{11.43}$ ${76.19}_{9.52}$ ${80.0}_{9.18}$ ${47.06}_{12.48}$ Effects of Act. ${82.61}_{8.08}$ ${66.67}_{12.6}$ ${85.71}_{9.71}$ ${77.27}_{9.14}$ ${100.0}_{0.0}$ ${60.0}_{13.09}$ ${80.0}_{10.69}$ ${87.5}_{8.54}$ Numerical RAC ${64.71}_{11.95}$ ${53.85}_{14.39}$ ${46.67}_{13.33}$ ${50.0}_{13.87}$ ${35.71}_{13.29}$ ${44.44}_{17.57}$ ${31.58}_{10.96}$ ${31.82}_{10.16}$ Composite ${78.57}_{6.41}$ ${67.5}_{7.5}$ ${68.42}_{7.64}$ ${70.73}_{7.19}$ ${66.67}_{7.65}$ ${53.85}_{8.09}$ ${60.98}_{7.71}$ ${78.38}_{6.86}$ Average ${78.95}_{3.55}$ ${73.45}_{4.17}$ ${68.87}_{4.52}$ ${77.44}_{3.64}$ ${73.21}_{4.2}$ ${64.04}_{4.51}$ ${65.89}_{4.19}$ ${70.69}_{4.24}$ 10 Fluent Tracking ${75.0}_{25.0}$ ${91.67}_{8.33}$ ${100.0}_{0.0}$ ${100.0}_{0.0}$ ${80.0}_{13.33}$ ${73.68}_{10.38}$ ${76.19}_{9.52}$ ${94.74}_{5.26}$ State Tracking ${86.67}_{9.09}$ ${63.64}_{15.21}$ ${83.33}_{11.24}$ ${73.33}_{11.82}$ ${63.64}_{15.21}$ ${61.54}_{14.04}$ ${81.25}_{10.08}$ ${87.5}_{12.5}$ Action Executability ${52.38}_{11.17}$ ${64.29}_{13.29}$ ${73.68}_{10.38}$ ${57.14}_{11.07}$ ${46.67}_{13.33}$ ${64.71}_{11.95}$ ${65.0}_{10.94}$ ${55.0}_{11.41}$ Effects of Act. ${68.0}_{9.52}$ ${66.67}_{10.54}$ ${78.57}_{11.38}$ ${69.57}_{9.81}$ ${78.57}_{11.38}$ ${30.77}_{13.32}$ ${61.54}_{14.04}$ ${66.67}_{10.54}$ Numerical RAC ${57.14}_{11.07}$ ${44.0}_{10.13}$ ${43.75}_{12.81}$ ${37.5}_{18.3}$ ${33.33}_{12.6}$ ${42.86}_{11.07}$ ${29.41}_{11.39}$ ${27.27}_{9.72}$ Composite ${71.74}_{6.71}$ ${78.05}_{6.54}$ ${74.42}_{6.73}$ ${77.55}_{6.02}$ ${55.56}_{8.4}$ ${54.35}_{7.43}$ ${64.29}_{7.48}$ ${71.43}_{7.75}$ Average ${67.42}_{4.09}$ ${67.74}_{4.21}$ ${74.36}_{4.05}$ ${73.91}_{3.75}$ ${57.43}_{4.94}$ ${55.04}_{4.4}$ ${63.57}_{4.25}$ ${64.8}_{4.29}$ 19 Fluent Tracking ${93.75}_{6.25}$ ${66.67}_{12.6}$ ${62.5}_{18.3}$ ${100.0}_{0.0}$ ${71.43}_{18.44}$ ${93.33}_{6.67}$ ${85.71}_{9.71}$ ${61.54}_{14.04}$ State Tracking ${91.67}_{5.76}$ ${81.25}_{10.08}$ ${86.67}_{9.09}$ ${70.0}_{15.28}$ ${87.5}_{8.54}$ ${58.33}_{14.86}$ ${78.57}_{11.38}$ ${83.33}_{16.67}$ Action Executability ${52.17}_{10.65}$ ${64.29}_{13.29}$ ${57.89}_{11.64}$ ${70.0}_{10.51}$ ${45.45}_{10.87}$ ${56.25}_{12.81}$ ${58.82}_{12.3}$ ${68.18}_{10.16}$ Effects of Act. ${72.73}_{9.72}$ ${80.95}_{8.78}$ ${85.71}_{9.71}$ ${70.0}_{10.51}$ ${76.92}_{12.16}$ ${41.67}_{14.86}$ ${66.67}_{11.43}$ ${76.47}_{10.6}$ Numerical RAC ${50.0}_{11.47}$ ${57.89}_{11.64}$ ${38.46}_{14.04}$ ${30.0}_{10.51}$ ${12.5}_{8.54}$ ${37.5}_{12.5}$ ${26.67}_{11.82}$ ${31.82}_{10.16}$ Composite ${72.92}_{6.48}$ ${60.47}_{7.54}$ ${72.09}_{6.92}$ ${79.55}_{6.15}$ ${53.85}_{8.09}$ ${55.1}_{7.18}$ ${67.44}_{7.23}$ ${86.21}_{6.52}$ Average ${71.9}_{3.65}$ ${67.19}_{4.17}$ ${68.75}_{4.4}$ ${70.31}_{4.05}$ ${54.87}_{4.7}$ ${56.67}_{4.54}$ ${64.46}_{4.37}$ ${66.97}_{4.53}$
Table 8: Performance across domains on GPT-4o on both the binary and free-answer subsets of the benchmark, evaluated without the ramifications constraints using the zero-shot setting. The results are categorized by the action-sequence lengths and question categories.
Table 8 shows the results of the binary questions across every domain in our benchmark.
C.4 Results by Fluents
Action Seq. Fluent Types GPT-4o Llama-8B-Inst Llama-70B-Inst Finetuned Llama-8B 1 Base Fluents ${81.25}_{5.63}$ ${65.22}_{7.02}$ ${74.55}_{5.87}$ ${96.3}_{2.57}$ Derived Fluents ${82.81}_{4.72}$ ${76.36}_{5.73}$ ${80.28}_{4.72}$ ${86.75}_{3.72}$ Self-Derived Fluents ${85.87}_{3.63}$ ${68.24}_{5.05}$ ${81.44}_{3.95}$ ${98.04}_{1.37}$ Static Properties ${85.19}_{6.84}$ ${72.73}_{7.75}$ ${80.0}_{6.32}$ ${89.47}_{4.06}$ 10 Base Fluents ${88.89}_{4.68}$ ${78.79}_{7.12}$ ${75.93}_{5.82}$ ${90.57}_{4.02}$ Derived Fluents ${75.31}_{4.79}$ ${55.38}_{6.17}$ ${72.73}_{5.08}$ ${79.38}_{4.11}$ Self-Derived Fluents ${77.66}_{4.3}$ ${74.29}_{5.22}$ ${69.39}_{4.66}$ ${95.15}_{2.12}$ Static Properties ${79.41}_{6.93}$ ${63.16}_{7.83}$ ${69.77}_{7.0}$ ${94.52}_{2.66}$ 19 Base Fluents ${90.24}_{4.63}$ ${63.64}_{8.37}$ ${76.19}_{6.57}$ ${83.61}_{4.74}$ Derived Fluents ${74.42}_{4.7}$ ${44.07}_{6.46}$ ${67.47}_{5.14}$ ${76.47}_{4.2}$ Self-Derived Fluents ${80.0}_{4.34}$ ${70.91}_{6.12}$ ${71.74}_{4.69}$ ${93.46}_{2.39}$ Static Properties ${82.76}_{7.01}$ ${60.71}_{9.23}$ ${67.5}_{7.41}$ ${96.88}_{2.17}$
Table 9: Performance comparison of GPT-4o, Llama-3.1-8B-Instruct, Llama-3.1-70B-Instruct, and fine-tuned Llama-3.1-8B on the binary subset (True/False) of the benchmark evaluated without the ramifications constraints using the zero-shot-CoT. The results are categorized up by the fluent types
Action Seq. Fluent Types GPT-4o Llama-8B-Inst Llama-70B-Inst Finetuned Llama-8B 1 Positive Fluents ${83.16}_{3.84}$ ${80.0}_{5.39}$ ${79.31}_{4.34}$ ${72.84}_{4.94}$ Negative Fluents ${77.78}_{4.62}$ ${60.34}_{6.42}$ ${70.77}_{5.64}$ ${77.53}_{4.42}$ Pos. and Neg. Fluents ${81.56}_{2.08}$ ${65.15}_{2.72}$ ${78.65}_{2.17}$ ${93.77}_{1.29}$ 10 Positive Fluents ${82.47}_{3.86}$ ${80.77}_{5.47}$ ${81.16}_{4.71}$ ${80.0}_{4.34}$ Negative Fluents ${66.27}_{5.19}$ ${53.66}_{7.79}$ ${69.23}_{5.72}$ ${72.29}_{4.91}$ Pos. and Neg. Fluents ${71.32}_{2.32}$ ${67.42}_{2.88}$ ${68.21}_{2.43}$ ${87.74}_{1.73}$ 19 Positive Fluents ${79.31}_{4.34}$ ${67.44}_{7.15}$ ${72.86}_{5.32}$ ${84.93}_{4.19}$ Negative Fluents ${71.08}_{4.98}$ ${58.14}_{7.52}$ ${67.53}_{5.34}$ ${75.64}_{4.86}$ Pos. and Neg. Fluents ${70.98}_{2.33}$ ${54.0}_{3.52}$ ${67.5}_{2.47}$ ${86.27}_{1.7}$
Table 10: Performance comparison of GPT-4o, Llama-3.1-8B-Instruct, Llama-3.1-70B-Instruct, and fine-tuned Llama-3.1-8B on the binary subset (True/False) of the benchmark evaluated without the ramifications constraints using the zero-shot-CoT. The results are categorized up by the fluent types
Tables 9, 10 show the results of the binary questions across fluent types in our benchmark.
Appendix D Data Verification
@@
Appendix E Fine-tuning Details
In this section, we describe the fine-tuning performed on the training split of ActionReasoningBench described in section 3.5. We fine-tuned Llama-3.1-8B separately for binary (true/false) and free answer questions, using 6 epochs for the former and 18 epochs for the latter. The AdamW optimizer was used, with a batch size of 4 and gradient accumulation steps set to 8 for both of the training setups. Due to the available compute resources, we were limited to a maximum context length of 4096 tokens. This leaves us roughly with 27k samples for binary answers and 14.4k samples for free answer. Tables 12, 13 and 14 show the statistics of the training set that we used to train the models.
@@ Appendix E Fine-tuning Details
Appendix F Free Answers Evaluation Details
@@
Table 15 shows the results of the free-answers using automatic metrics based on semantic similarity, F1-BERTScore (Zhang et al., 2019).
Action Seq. Ques Categories GPT-4o Llama-8B-Inst Llama-70B-Inst Finetuned Llama-8B 1 Fluent Tracking ${71.76}_{2.11}$ ${48.52}_{2.84}$ ${55.21}_{2.85}$ ${90.22}_{0.95}$ State Tracking ${68.24}_{1.98}$ ${54.85}_{3.18}$ ${61.07}_{2.59}$ ${82.06}_{1.82}$ Action Executability ${56.93}_{3.36}$ ${49.49}_{2.80}$ ${57.25}_{3.77}$ ${75.24}_{2.70}$ Effects of Actions ${66.43}_{1.94}$ ${49.82}_{3.53}$ ${59.48}_{2.67}$ ${82.12}_{2.07}$ Numerical RAC ${87.07}_{1.61}$ ${39.40}_{5.05}$ ${61.94}_{5.30}$ ${89.52}_{1.36}$ Composite Questions ${55.03}_{1.65}$ ${40.59}_{1.52}$ ${40.62}_{1.77}$ ${82.58}_{1.17}$ Average ${67.55}_{2.10}$ ${47.11}_{3.15}$ ${55.92}_{3.15}$ ${83.62}_{1.67}$ 10 Fluent Tracking ${66.46}_{2.29}$ ${51.24}_{2.98}$ ${56.89}_{3.20}$ ${85.40}_{1.74}$ State Tracking ${70.09}_{1.26}$ ${60.43}_{2.52}$ ${64.10}_{2.23}$ ${83.78}_{1.49}$ Action Executability ${61.31}_{3.51}$ ${54.14}_{2.79}$ ${68.28}_{2.78}$ ${79.04}_{2.40}$ Effects of Actions ${66.19}_{1.34}$ ${52.99}_{3.09}$ ${62.05}_{2.16}$ ${82.03}_{1.78}$ Numerical RAC ${82.38}_{1.82}$ ${29.14}_{4.13}$ ${50.39}_{5.46}$ ${85.20}_{2.62}$ Composite Questions ${54.57}_{1.68}$ ${45.67}_{1.59}$ ${54.41}_{1.39}$ ${79.83}_{1.56}$ Average ${66.83}_{1.98}$ ${48.93}_{2.85}$ ${59.35}_{2.87}$ ${82.47}_{1.93}$ 19 Fluent Tracking ${60.78}_{3.19}$ ${47.69}_{2.71}$ ${56.35}_{3.12}$ ${87.79}_{1.39}$ State Tracking ${70.33}_{1.00}$ ${58.06}_{2.77}$ ${62.06}_{2.47}$ ${83.51}_{1.41}$ Action Executability ${61.72}_{3.41}$ ${49.40}_{2.93}$ ${56.90}_{3.48}$ ${78.84}_{2.40}$ Effects of Actions ${65.71}_{1.99}$ ${52.53}_{3.51}$ ${62.78}_{3.22}$ ${84.90}_{1.20}$ Numerical RAC ${81.30}_{2.29}$ ${32.55}_{4.31}$ ${53.79}_{5.36}$ ${80.75}_{3.62}$ Composite Questions ${55.48}_{1.55}$ ${50.87}_{1.40}$ ${52.75}_{1.63}$ ${77.10}_{1.76}$ Average ${65.88}_{2.23}$ ${48.51}_{1.95}$ ${57.43}_{3.21}$ ${82.14}_{1.96}$
Table 15: F1-BERTScore of GPT-4o, Llama-3.1-8B-Instruct, Llama-3.1-70B-Instruct, and fine-tuned Llama-3.1-8B on the free-answer subset of the benchmark, evaluated without the ramifications constraints using the zero-shot-CoT. The results are categorized up by the action-sequence lengths and question categories.
Below we share some examples where F1-BertScore fails to give an accurate score.
Low F1-BERTScore but the response is correct [GROUND TRUTH]
robot1’s rgripper1 is available
[RESPONSE]
rgripper1 is free.
[BERT SCORE]
0.71
[Llama-3.1-70B-Intruct Evaluation]
The LLM response and the ground truth response are semantically the same. The term ”free” in the LLM response is equivalent to ”available” in the ground truth response, both indicating that the rgripper1 is not being used. True.
[ANALYSIS]
The response is a simple paraphrase of the ground truth which is easily explained by the LLM evaluation.
High F1-BERTScore but the response is incorrect [GROUND TRUTH]
loc_x0_y1 is not connected to loc_x2_y4
loc_x0_y3 is not connected to loc_x2_y4
loc_x1_y0 and loc_x2_y4 are not connected
loc_x1_y1 is not connected to loc_x2_y4
there is no connection between loc_x3_y0 and loc_x2_y4
there is no connection between loc_x3_y1 and loc_x2_y4
there is no connection between loc_x3_y2 and loc_x2_y4
there is no connection between loc_x3_y3 and loc_x2_y4
[RESPONSE]
- loc_x2_y4 is not visited.
- loc_x2_y4 is connected to loc_x1_y4.
- loc_x2_y4 is connected to loc_x3_y4.
- loc_x2_y4 is not connected to loc_x2_y3.
- loc_x2_y4 is not connected to loc_x2_y5.
- loc_x2_y4 is not connected to loc_x1_y2.
- loc_x2_y4 is not connected to loc_x1_y3.
- loc_x2_y4 is not connected to loc_x1_y5.
[BERT SCORE]
0.89
[Llama-3.1-70B-Intruct Evaluation]
The LLM response and the ground truth response are not semantically the same. The LLM response states that loc_x2_y4 is not connected to loc_x2_y3, but the ground truth response states that loc_x2_y4 and loc_x2_y3 are connected. Additionally, the LLM response states that loc_x2_y4 is not connected to loc_x2_y5, but the ground truth response states that loc_x2_y4 and loc_x2_y5 are connected. False.
[ANALYSIS]
The response contains a lot of incorrect responses. Moreover, the response fails to mention the location of the robot. The Llama-3.1-70B-Instruct evaluation also fails to mention the lack of the robot but correctly identifies some incorrect connections.
Appendix G Domains in ActionReasoningBench
@@ G.8 VisitAll
Appendix H Planning Description and Tools
@@
H.1 Planning Domain Definition Language (PDDL)